OpenStudy (anonymous):

Show that the equation represents a circle by rewriting it in standard form. x^2+y^2-(1/4x)+(1/4y)=(1/32)

OpenStudy (anonymous):

Complete the square for both x and y terms.. (x^2-(1/4)x+(1/64)) + (y^2+(1/4)y+(1/64))=(1/32)+(1/64)+(1/64) (x-(1/8))^2+(y+(1/8))^2 = 1/16 Center = (1/8,-1/8) Radius = 1/4